# On the Closed-Form of Flow Matching: Generalization Does Not Arise from Target Stochasticity

### v1

@@ 3.2 Failure to learn the optimal velocity field
Set up of Figure 3. To build Figure 3, we subsampled the CIFAR-10 dataset from $10$ to $10^{4}$ samples. For each size, we trained a flow matching model using a standard $34$ million-parameter U-Net (see Appendix C for details). Following Kadkhodaie et al. (2024), the number of parameters of the network $u_{\theta}$ remains fixed across dataset sizes. Importantly, the optimal velocity field $\hat{u}^{\star}$ itself depends on the dataset size: as the number of samples increases, the complexity of $\hat{u}^{\star}$ also grows. Thus, we expect the network $u_{\theta}$ to accurately approximate the optimal velocity field $\hat{u}^{\star}$ for smaller dataset sizes.
@@
Comments on Figure 3. The leftmost plot shows the average training error
$\mathbb{E}_{\begin{subarray}{c}x_{0}\sim p_{0}\\ x_{1}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left\|u_{\theta}(x_{t},t)-\hat{u}^{\star}(x_{t},t)\right\|^{2},\quad\text{where}\quad x_{t}:=(1-t)x_{0}+tx_{1}\enspace,$
between the learned velocity $u_{\theta}$ and the optimal empirical velocity field $\hat{u}^{\star}$, evaluated across multiple time values $t$ and dataset sizes. With only $10$ samples (darkest curve), the network $u_{\theta}$ closely approximates $\hat{u}^{\star}$. As the dataset size increases, the complexity of $\hat{u}^{\star}$ grows, and the approximation by $u_{\theta}$ becomes less accurate. In particular, the approximation fails at two specific time intervals: around $t\approx 0.15$ and near $t=1$. The failure near $t=1$ is expected, as $\hat{u}^{\star}$ becomes non-Lipschitz at $t=1$. Interestingly, the early-time failure at $t\approx 0.15$ corresponds to the regime where $\hat{u}^{\star}$ and $u^{\mathrm{cond}}$ start to correlate (see Figure 1(b) in Section 3.1). The middle plot of Figure 3 reports the FID-10k, computed on the test set in the DINOv2 embedding space (Oquab et al., 2024), for various dataset sizes. For a small dataset (e.g., #$\#\text{samples}=10$), $u_{\theta}$ approximates $\hat{u}^{\star}$ well but does not generalize – the test FID exceeds $10^{3}$. As the dataset size increases ($500\leq\#\text{samples}\leq 3000$), the approximation $u_{\theta}$ becomes less accurate. Despite this, the model achieves lower FID scores on the test set but still memorizes the training data. The rightmost plot of Figure 3 illustrates this memorization by showing the average distance between each generated sample and its nearest neighbor in the training set. For larger datasets (#$\#\text{samples}\geq 3000$), this distance increases substantially, indicating that the model generalizes better. Overall, Figure 3 also suggests that the FID metric can be misleading, even when computed on the test set. For example, the model trained with 1000 samples has a low test FID but memorizes training examples.
Figure 3 confirms that generalization arises when the network $u_{\theta}$ fails to estimate the optimal velocity field $\hat{u}^{\star}$, and that this failure occurs at two specific time intervals. In Section 3.3, we investigate which of these two intervals is responsible for driving generalization.
@@ 3.3 When does generalization arise?
Set up of Figure 4. We first learn a velocity field $u_{\theta}$ using standard conditional flow matching (see Appendix C), then we construct a hybrid model: we define a piecewise trajectory where the flow is governed by the optimal velocity field $\hat{u}^{\star}$ for times $t\in[0,\tau]$, and by the learned velocity field $u_{\theta}$ for times $t\in[\tau,1]$, for a given threshold parameter $\tau\in[0,1]$. For the extreme case $\tau=1$, the full trajectory follows $\hat{u}^{\star}$, and samples exactly match training data points. Conversely, when $\tau=0$, the entire trajectory is governed by $u_{\theta}$, yielding novel samples. Intermediate values of $\tau$ produce a mixture of both behaviors, which we interpret as reflecting varying degrees of generalization. To assess generalization, we measure the distance of generated samples to the dataset using the LPIPS metric (Zhang et al., 2018), which computes the feature distance between two images via some pretrained classification network. We define the distance of a generated sample $x$ to a dataset $\mathcal{D}=\{x^{(1)},\ldots,x^{(n)}\}$ as $\mathrm{dist}(x,\mathcal{D})=\min_{x^{(i)}\in\mathcal{D}}\mathrm{LPIPS}(x,x^{(i)})$.
@@
Figure 4: Generalization occurs at small times on CIFAR-10 (left) and CelebA64646464 (right). Top: Generalization (distance between generated samples and training data) of hybrid models that follow $\hat{u}^{\star}$ on $[0,\tau]$, then $u_{\theta}$ on $[\tau,1]$. The four colored curves correspond to four specific $x_{0}$, the black dashed curve is the mean distance over the 256 generated images. Bottom: visualization of generated images for the four different starting noises and various values of $\tau$ (the background color matching the curve in the top figure). Following $\hat{u}^{\star}$ until $\tau\geq 0.3$ yields a model that is not able to generalize.
@@
We fix a random batch of 256 pure noise images from $p_{0}$. Then, for various threshold values $\tau$, we generate 256 images with the hybrid model, always starting from this batch. Finally, we measure the creativity of the hybrid model as the mean of the aforementioned LPIPS distances between the 256 generated samples and the dataset. Figure 4 (top row) displays these distances as $\tau$ varies, on the CIFAR-10 (left) and CelebA - $64\times 64$ (right) datasets. For $\tau\leq 0.2$, the hybrid model remains as creative as $u_{\theta}$, despite following $\hat{u}^{\star}$ in the first steps. For $\tau>0.2$, the LPIPS distance starts dropping. On the displayed generated samples (bottom rows), we in fact see that as soon as $\tau\geq 0.4$, the sample generated by the hybrid model is almost the same as the one obtained with $\hat{u}^{\star}$ $(\tau=1)$. This means that the final image is already determined at $t=0.4$ and despite the generalization capacity of $u_{\theta}$, following it only after $t\geq 0.4$ is not to enough to create a new image. Hence, generalization occurs early and cannot be explained by the failure to correctly approximate $u^{\star}$ at large $t$.
@@
Although we have shown that the stochastic phase was limited to small values of $t$ in real-data settings, we have not yet definitively ruled it out as the cause of generalization. In the following section, we introduce a learning procedure designed to address this question directly.
@@ 4 Learning with the closed-form formula
In this section, to discard the impact of stochastic target on the generalization, we propose to directly regress against the closed-form formula (5).
@@ 4.1 Empirical flow matching
Regressing against the closed-form $\hat{u}^{\star}$ at a point $(x_{t},t)$ requires evaluating it, i.e., computing a weighted sum of the conditional velocity fields over all the $n$ training points $x^{(i)}$. For a dataset of $n$ samples of size $d$, and a batch of size $|\mathcal{B}|$, computing the weights of the exact closed-form formula $\hat{u}^{\star}(x,t)$ of flow matching requires $\mathcal{O}(n\times|\mathcal{B}|\times d)$. These computations are prohibitive since they must be performed for each batch. One natural idea is to rely on a Monte Carlo approximation of the closed-form formula, using $M$ samples $b^{(1)},\dots,b^{(M)}$ to estimate the empirical mean:
$\mathcal{L}_{\mathrm{EFM}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}([0,1])\\ x_{0}\sim p_{0}\\ x_{1}\sim\hat{p}_{\mathrm{data}}\\ x_{t}=(1-t)x_{0}+tx_{1}\\ b^{(1)}:=x_{1}\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\enspace,$ (7)
with
$\hat{u}_{M}^{\star}(x_{t},t)$ $=\sum_{j=1}^{M}\frac{b^{(j)}-x_{t}}{1-t}\cdot\left[\mathrm{softmax}\left(\left(-\frac{\lVert{x_{t}-tb^{(k)}}\rVert^{2}}{2(1-t)^{2}}\right)_{k=1,\dots,M}\right)\right]_{j}\enspace.$ (8)
The formulation in Equation 7 may appear naive at first glance. Still, it hinges on a crucial trick: the Monte Carlo estimate is computed using a batch that systematically includes the point $x_{1}$, that generated the current $x_{t}$. If instead $b^{(1)}$ were sampled independently from $\hat{p}_{\mathrm{data}}$, this could introduce a sampling bias (see Ryzhakov et al. 2024, Appendix B, and the corresponding OpenReview comments [^6] https://openreview.net/forum?id=XYDMAckWMa for an in-depth discussion). Proposition 2 shows that the estimate $\hat{u}_{M}^{\star}$ is unbiased and has lower variance than the standard conditional flow matching target.
@@
    The minimizer of Equation 7 writes, for all $(x,t)$
    $\mathbb{E}_{b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x,t)\right]\enspace.$ (9)
@@
    In addition, for all $(x,t)$, the minimizer of Equation 7 equals the optimal velocity field, i.e.,
    $\mathbb{E}_{b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x,t)\right]=\hat{u}^{\star}(x,t)\enspace.$ (10)
@@
    $\mathrm{Var}_{b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)\;;\;b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\left[\hat{u}_{M}^{\star}(x,t)\right]\leq\mathrm{Var}_{b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)}\left[u^{\mathrm{cond}}(x,b^{(1)},t)\right].$ (11)
The proof of Proposition 2 is provided in Section B.3. The estimator $\hat{u}_{M}^{\star}$ of the optimal field $\hat{u}^{\star}$ is closely related to self-normalized importance sampling (see Section B.2 and Owen 2013, Chap. 9.2), as well as to Rao-Blackwellized estimators (Casella and Robert, 1996; Cardoso et al., 2022). As discussed in Ryzhakov et al. (2024), self-normalized importance sampling estimators of $\hat{u}^{\star}$ are generally biased, in the sense that: $\mathbb{E}_{b^{(1)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\hat{u}_{M}^{\star}(x_{t},t)\neq\hat{u}^{\star}(x_{t},t)\enspace.$ A key insight is that our estimator includes $b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot\mid x_{t},t)$, which leads to the main result of Proposition 2. In Section 4.2, we demonstrate that Algorithm 2, designed to solve Equation 7, yields consistent improvements on high-dimensional datasets such as CIFAR-10 and CelebA. Additional details on the unbiasedness of $\mathcal{L}_{\mathrm{EFM}}$ can be found in the supplementary material (Appendix B). From a computational perspective, despite requiring $M$ additional samples, Algorithm 2 remains significantly more efficient than increasing the batch size by a factor of $M$: the $M$ samples are merely averaged (with weights), while the backpropagation remains identical to that of Algorithm 1.
@@
      ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u^{\mathrm{cond}}(x_{t},t)=x_{1}-x_{0}}$
@@ 4.2 Experiments
We now learn with empirical flow matching (EFM, Equation 7 and Algorithm 2) in practical high-dimensional settings. Our goal with this empirical investigation is first to observe if regressing against a more deterministic target leads to performance improvement/degradation.
@@
Datasets and Models. We perform experiments on the image datasets CIFAR-10 (Krizhevsky and Hinton, 2009) and CelebA $64\times 64$ (Liu et al., 2015). For the experiments, we compare vanilla conditional flow matching (Lipman et al., 2023; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023), optimal transport flow matching (Pooladian et al., 2023; Tong et al., 2024), and the empirical flow matching in Algorithm 2, for multiple numbers of samples $M$ to estimate the empirical mean. Training details are in Appendix C.
@@
Metrics. To assess generalization performance, we use the standard Fréchet Inception Distance (Heusel et al., 2017) with Inception-V3 (Szegedy et al., 2016) but we also follow the recommendation of Stein et al. (2023) using the DINOv2 embedding (Oquab et al., 2023), which is known to lead to a less biased evaluation. Defining and quantifying the generalization ability of generative models is overall a challenging task: train and test FID are known to be imperfect (Stein et al., 2023; Jiralerspong et al., 2023; Parmar et al., 2022) yet no superior competitor has emerged.
@@
Figure 5: FID computed on the training set ($50$k) and the test set ($10$k) using multiple embeddings, Inception and DINOv2. Regressing against a more deterministic target (EFM - $128$, $256$, $1000$) does not yield performance decreases. On the contrary, the more deterministic the target, the better the performance.
@@
Comments on Figure 5. Figure 5 compares vanilla flow matching, OTCFM, and the empirical flow matching (EFM, Algorithm 2) approaches using various numbers of samples to estimate the empirical mean, $M\in\{128,256,1000\}$. First, we observe that learning with a more deterministic target does not degrade either training or testing performance, across both types of embeddings. On the contrary, we consistently observe modest but steady improvements as stochasticity is reduced. For both CIFAR-10 and CelebA, increasing the number of samples $M$ used to compute the empirical mean—i.e., , making the targets less stochastic—leads to more stable improvements. It is worth noting that Algorithm 2 has a computational complexity of $\mathcal{O}(M\times|\mathcal{B}|\times d)$, where $|\mathcal{B}|$ is the batch size, $M$ is the number of samples used to estimate the empirical mean, and $d$ is the sample dimension. In our experiments, choosing $M=|\mathcal{B}|=128$ yielded a modest time overhead.

### v2

@@ 3.2 Failure to learn the optimal velocity field
Set up of Figure˜2. To build Figure˜2, we subsampled the CIFAR-10 dataset from $10$ to $10^{4}$ samples. For each size, we trained a flow matching model using a standard $34$ million-parameter U-Net (see Appendix˜D for details). Following kadkhodaie2024generalization, the number of parameters of the network $u_{\theta}$ remains fixed across dataset sizes. Importantly, the optimal velocity field $\hat{u}^{\star}$ itself depends on the dataset size: as the number of samples increases, the complexity of $\hat{u}^{\star}$ also grows. Thus, we expect the network $u_{\theta}$ to accurately approximate the optimal velocity field $\hat{u}^{\star}$ for smaller dataset sizes.
@@
Comments on Figure˜2. The leftmost plot shows the average training error
$\mathbb{E}_{\begin{subarray}{c}x_{0}\sim p_{0}\\ x_{1}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left\|u_{\theta}(x_{t},t)-\hat{u}^{\star}(x_{t},t)\right\|^{2},\quad\text{where}\quad x_{t}:=(1-t)x_{0}+tx_{1}\kern 5.0pt,$
between the learned velocity $u_{\theta}$ and the optimal empirical velocity field $\hat{u}^{\star}$, evaluated across multiple time values $t$ and dataset sizes. With only $10$ samples (darkest curve), the network $u_{\theta}$ closely approximates $\hat{u}^{\star}$. As the dataset size increases, the complexity of $\hat{u}^{\star}$ grows, and the approximation by $u_{\theta}$ becomes less accurate. In particular, the approximation fails at two specific time intervals: around $t\approx 0.15$ and near $t=1$. The failure near $t=1$ is expected, as $\hat{u}^{\star}$ becomes non-Lipschitz at $t=1$. Interestingly, the early-time failure at $t\approx 0.15$ corresponds to the regime where $\hat{u}^{\star}$ and $u^{\mathrm{cond}}$ start to correlate (see Figure˜1(a) in Section˜3.1). The middle plot of Figure˜2 reports the FID-10k, computed on the test set in the DINOv2 embedding space (dinov2), for various dataset sizes. For a small dataset (e.g., $\text{samples}=10$), $u_{\theta}$ approximates $\hat{u}^{\star}$ well but does not generalize – the test FID exceeds $10^{3}$. As the dataset size increases ($1000\leq\#\text{samples}\leq 3000$), the approximation $u_{\theta}$ becomes less accurate. Despite this, the model achieves lower FID scores on the test set but still memorizes the training data. The rightmost plot of Figure˜2 illustrates this memorization by showing the average distance between each generated sample and its nearest neighbor in the training set. For larger datasets ($\text{samples}\geq 3000$), this distance increases substantially, indicating that the model generalizes better. Overall, Figure˜2 also suggests that the FID metric can be misleading, even when computed on the test set. For example, the model trained with 1000 samples has a low test FID but memorizes training examples.
Figure˜2 confirms that generalization arises when the network $u_{\theta}$ fails to estimate the optimal velocity field $\hat{u}^{\star}$, and that this failure occurs at two specific time intervals. In Section˜3.3, we investigate which of these two intervals is responsible for driving generalization.
@@ 3.3 When does generalization arise?
Figure 3: Generalization occurs at small times on CIFAR-10 (left) and CelebA $64$ (right). Top: Generalization (distance between generated samples and training data) of hybrid models that follow $\hat{u}^{\star}$ on $0,\tau]$, then $u_{\theta}$ on $\tau,1]$. The four colored curves correspond to four specific $x_{0}$, the black dashed curve is the mean distance over the 256 generated images. Bottom: visualization of generated images for the four different starting noises and various values of $\tau$ (the background color matching the curve in the top figure). Following $\hat{u}^{\star}$ until $\tau\geq 0.3$ yields a model that is not able to generalize.
Set up of Figure˜3. We first learn a velocity field $u_{\theta}$ using standard conditional flow matching (see Appendix˜D), then we construct a hybrid model: we define a piecewise trajectory where the flow is governed by the optimal velocity field $\hat{u}^{\star}$ for times $t\in[0,\tau]$, and by the learned velocity field $u_{\theta}$ for times $t\in[\tau,1]$, for a given threshold parameter $\tau\in[0,1]$. For the extreme case $\tau=1$, the full trajectory follows $\hat{u}^{\star}$, and samples exactly match training data points. Conversely, when $\tau=0$, the entire trajectory is governed by $u_{\theta}$, yielding novel samples. Intermediate values of $\tau$ produce a mixture of both behaviors, which we interpret as reflecting varying degrees of generalization. To assess generalization, we measure the distance of generated samples to the dataset using the LPIPS metric (zhang2018unreasonable), which computes the feature distance between two images via some pretrained classification network. We define the distance of a generated sample $x$ to a dataset $\mathcal{D}=\{x^{(1)},\ldots,x^{(n)}\}$ as $\mathrm{dist}(x,\mathcal{D})=\min_{x^{(i)}\in\mathcal{D}}\mathrm{LPIPS}(x,x^{(i)})$. We fix a random batch of 256 pure noise images from $p_{0}$. Then, for various threshold values $\tau$, we generate 256 images with the hybrid model, always starting from this batch. Finally, we measure the creativity of the hybrid model as the mean of the aforementioned LPIPS distances between the 256 generated samples and the dataset.
@@
Comments on Figure˜3. The top row displays the LPIPS distances as $\tau$ varies, on the CIFAR-10 (left) and CelebA - $64\times 64$ (right) datasets. For $\tau\leq 0.2$, the hybrid model remains as creative as $u_{\theta}$, despite following $\hat{u}^{\star}$ in the first steps. For $\tau>0.2$, the LPIPS distance starts dropping. On the displayed generated samples (bottom rows), we in fact see that as soon as $\tau\geq 0.4$, the sample generated by the hybrid model is almost the same as the one obtained with $\hat{u}^{\star}$ $(\tau=1)$. This means the final image is already determined at $t=0.4$, and despite the generalization capacity of the learned velocity field $u_{\theta}$, following it only after $t\geq 0.4$ is not enough to create a new image: generalization occurs early and cannot fully be explained by the failure to correctly approximate $u^{\star}$ at large $t$.
@@
Although we have shown that the stochastic phase was limited to small values of $t$ in real-data settings, we have not yet definitively ruled it out as the cause of generalization. In the following Section˜4, we introduce a learning procedure designed to address this question directly.
@@ 4 Learning with the closed-form formula
In this section, in order to discard the impact of stochastic target on the generalization, we propose to directly regress against the closed-form formula in Equation˜6.
@@ 4.1 Empirical flow matching
Regressing against the closed-form $\hat{u}^{\star}$, defined in Equation˜6, at a point $(x_{t},t)$ requires computing a weighted sum of the conditional velocity fields over all the $n$ training points $x^{(i)}$. For a dataset of $n$ samples of size $d$, and a batch of size $|\mathcal{B}|$, computing the weights of the exact closed-form formula $\hat{u}^{\star}(x,t)$ of flow matching requires $\mathcal{O}(n\times|\mathcal{B}|\times d)$. These computations are prohibitive since they must be performed for each batch. One natural idea is to estimate the closed-form formula $\hat{u}^{\star}$ (Equation˜6), by a Monte Carlo approximation (Equation˜8), using $M\leq n$ samples $b^{(1)},\dots,b^{(M)}$:
$\mathcal{L}_{\mathrm{EFM}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}x_{0}\sim p_{0}\\ x_{1}\sim\hat{p}_{\mathrm{data}}\\ t\sim\mathcal{U}([0,1])\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\|u_{\theta}(x_{t},t)-\hat{u}_{M}^{\star}(x_{t},t)\|^{2}\kern 5.0pt,$ (7)
with $x_{t}=(1-t)x_{0}+tx_{1}$, $b^{(1)}:=x_{1}$, and
$\hat{u}_{M}^{\star}(x,t)$ $=\sum_{j=1}^{M}\lambda(x,t)\frac{b^{(j)}-x}{1-t}\kern 5.0pt,\kern 5.0pt\lambda(x,t)=\mathrm{softmax}\left(\left(-\frac{\|x-tx^{(l)}\|^{2}}{2(1-t)^{2}}\right)_{l=1,\ldots,n}\right)\kern 5.0pt.$ (8)
The formulation in Equation˜7 may appear naive at first glance. Still, it hinges on a crucial trick: the Monte Carlo estimate is computed using a batch that systematically includes the point $x_{1}$, that generated the current $x_{t}$. If instead $b^{(1)}$ were sampled independently from $\hat{p}_{\mathrm{data}}$, this could introduce a sampling bias (see Ryzhakov2024explicit, Appendix˜B, and the corresponding OpenReview comments [^2] https://openreview.net/forum?id=XYDMAckWMa for an in-depth discussion). Proposition˜2 shows that the estimate $\hat{u}_{M}^{\star}$ is unbiased and has lower variance than the standard conditional flow matching target.
@@
    The minimizer of Equation˜7 writes, for all $(x,t)$
    $\mathbb{E}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left[\hat{u}_{M}^{\star}(x,t)\right]\kern 5.0pt.$ (9)
@@
    In addition, for all $(x,t)$, the minimizer of Equation˜7 equals the optimal velocity field, i.e.,
    $\mathbb{E}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left[\hat{u}_{M}^{\star}(x,t)\right]=\hat{u}^{\star}(x,t)\kern 5.0pt.$ (10)
@@
    $\mathrm{Var}_{\begin{subarray}{c}b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)\\ b^{(2)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}\end{subarray}}\left[\hat{u}_{M}^{\star}(x,t)\right]\leq\mathrm{Var}_{b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot|x,t)}\left[u^{\mathrm{cond}}(x,b^{(1)},t)\right].$ (11)
The proof of Proposition˜2 is provided in Section˜B.3. The estimator $\hat{u}_{M}^{\star}$ of the optimal field $\hat{u}^{\star}$ is closely related to self-normalized importance sampling (see Section˜B.2 and Owen2013montecarlobook), as well as to Rao-Blackwellized estimators (casella1996raoblaclwell; cardoso2022biasreduced). As discussed in Ryzhakov2024explicit, self-normalized importance sampling estimators of $\hat{u}^{\star}$ are generally biased, in the sense that: $\mathbb{E}_{b^{(1)},\dots,b^{(M)}\sim\hat{p}_{\mathrm{data}}}\hat{u}_{M}^{\star}(x_{t},t)\neq\hat{u}^{\star}(x_{t},t)\kern 5.0pt.$ A key insight is that our estimator includes $b^{(1)}\sim\hat{p}_{\mathrm{data}}(\cdot\mid x_{t},t)$, which leads to the main result of Proposition˜2. In Section˜4.2, we demonstrate that Algorithm˜2, designed to solve Equation˜7, yields consistent improvements on high-dimensional datasets such as CIFAR-10 and CelebA. Additional details on the unbiasedness of $\mathcal{L}_{\mathrm{EFM}}$ can be found in the supplementary material (Appendix˜B). From a computational perspective, despite requiring $M$ additional samples, Algorithm˜2 remains significantly more efficient than increasing the batch size by a factor of $M$: the $M$ samples are merely averaged (with weights), while the backpropagation remains identical to that of Algorithm˜1.
@@
 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}u^{\mathrm{cond}}(x_{t},t)=\dfrac{x_{1}-x_{t}}{1-t}=x_{1}-x_{0}$}
@@ 4.2 Experiments
We now learn with empirical flow matching (EFM, Equation˜7 and Algorithm˜2) in practical high-dimensional settings. Our goal with this empirical investigation is first to observe if regressing against a more deterministic target leads to performance improvement/degradation.
@@
Datasets and Models. We perform experiments on the image datasets CIFAR-10 (Krizhevsky2009) and CelebA $64\times 64$ (celeba). For the experiments, we compare vanilla conditional flow matching (lipman2023flow; liu2023rectifiedflow; albergo2023stochasticinterpolant), optimal transport flow matching (pooladian2023multisample; tong2023improving), and the empirical flow matching in Algorithm˜2, for multiple numbers of samples $M$ to estimate the empirical mean. Training details are in Appendix˜D.
@@
Metrics. To assess generalization performance, we use the standard Fréchet Inception Distance (heusel2017gans) with Inception-V3 (szegedy2016rethinking) but we also follow the recommendation of stein2023evalgenmodels using the DINOv2 embedding (oquab2023dinov2), which is known to a more expressive and discriminative embedding, that leads to a less biased evaluation. We also measure the FID between the generated and the train and test sets, rather than only on the training set, as is often done in generative modeling benchmarks. The train FID is computed between 5050k generated images and the 5050k images from the training set. The test FID is computed between the same 5050k generated images and the 1010k images from the test set. On Figure˜2, we also displayed a memorization metric that would detect a pure copy of the training set. Overall, defining and quantifying the generalization ability of generative models is overall a challenging task: train and test FID are known to be imperfect (stein2023evalgenmodels; Jiralerspong2023fld; parmar2021cleanfid), yet no superior competitor has emerged.
@@
Figure 4: FID computed on the training set (5050k) and the test set (1010k) using multiple embeddings, Inception and DINOv2. Regressing against a more deterministic target (EFM - $128$, $256$, $1000$) does not yield performance decreases. On the contrary, the more deterministic the target, the better the performance.
@@
Comments on Figure˜4. Figure˜4 compares vanilla flow matching, OTCFM, and the empirical flow matching (EFM, Algorithm˜2) approaches using various numbers of samples to estimate the empirical mean, $M\in\{128,256,1000\}$. First, we observe that learning with a more deterministic target does not degrade either training or testing performance, across both types of embeddings. On the contrary, we consistently observe modest but steady improvements as stochasticity is reduced. For both CIFAR-10 and CelebA, increasing the number of samples $M$ used to compute the empirical mean—i.e., , making the targets less stochastic—leads to more stable improvements. It is worth noting that Algorithm˜2 has a computational complexity of $\mathcal{O}(M\times|\mathcal{B}|\times d)$, where $|\mathcal{B}|$ is the batch size, $M$ is the number of samples used to estimate the empirical mean, and $d$ is the sample dimension. In our experiments, choosing $M=|\mathcal{B}|=128$ yielded a modest time overhead. For empirical flow matching, we experimented with several values beyond $M=1000$ (e.g., $M=2000$, $M=5000$). The results were nearly identical to those obtained with $M=1000$, with curves being visually indistinguishable. Therefore, we chose not to report results for $M\geq 1000$.